OpenStudy (anonymous):

Freddie is at chess practice waiting on his opponent's next move. He notices that the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle.

OpenStudy (anonymous):

Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 3π inches? l = r(theta) 3pi = 4(theta) 3pi / 4 = theta Part 4: What is the coordinate point associated with this radian measure? \[(\sqrt[-2]{2},\sqrt[2]{2})\] Please help @oldrin.bataku @satellite73

OpenStudy (anonymous):

Just need someone to check answers

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Part 3 is correct. Part 4 depends on the interpretation of the movement of a chess clock. It's been a long time since I used one. I don't know if the hand turns clockwise (logical) or counter-clockwise ("like a unit circle"). Your answer is correct if the hand moves counter-clockwise.

OpenStudy (anonymous):

would \[(\sqrt[2]{2},\sqrt[-2]{2})\] change it?

OpenStudy (mathmate):

As I said, your answer is correct if the hand of the clock moves counter-clockwise ("135 degrees like a unit circle"). I do not know if a chess-clock moves counter-clockwise or clockwise. It is an interpretation of the question.

OpenStudy (mathmate):

At this link, the chess clock is shown to move clockwise. There is an ambiguity with the question. https://www.youtube.com/watch?v=14PmKfw_GSo